### v1

# Easy-access online social media metrics can effectively identify misinformation sharing users

Abstract
Misinformation poses a significant challenge studied extensively by researchers, yet acquiring data to identify primary sharers is costly and challenging. To address this, we propose a low-barrier approach to differentiate social media users who are more likely to share misinformation from those who are less likely. Leveraging insights from previous studies, we demonstrate that easy-access online social network metrics—average daily tweet count, and account age—can be leveraged to help identify potential low factuality content spreaders on X (previously known as Twitter). We find that higher tweet frequency is positively associated with low factuality in shared content, while account age is negatively associated with it. We also find that some of the effects, namely the effect of the number of accounts followed and the number of tweets produced, differ depending on the number of followers a user has. Our findings show that relying on these easy-access social network metrics could serve as a low-barrier approach for initial identification of users who are more likely to spread misinformation, and therefore contribute to combating misinformation effectively on social media platforms.
keywords:
Misinformation, social media, X, social considerations, cognitive load, digital literacy
Introduction
The problem of the global spread of misinformation has been recognised as one of the most pressing issues in today’s societies, prompting many researchers to dedicate efforts to understanding the issue as well as to developing potential solutions for tackling it [1]. Researchers have focused on a variety of areas to understand misinformation, examining the phenomenon from the perspectives of the misinformation producers, consumers, the messages themselves, and the contexts in which the messages are embedded [2]. While bots may play a significant role in spreading misinformation, the role of human users of sharing it remains a key element in the diffusion process [3]. Furthermore, as the spreading of misinformation carries the risk of subjecting people to manipulation and might influence behaviour in undesirable ways [4, 5], understanding the factors that affect individuals’ likelihood of sharing misinformation is essential.
The reasons behind people’s tendency to share misinformation have been widely studied, and findings have shown that in many cases the decision of users to share misinformation does not stem from malicious intent, but it is explained by a failure to consider the accuracy of the piece of news shared [6]. Failure to consider accuracy when sharing content has been shown to have two main sources: lack of attention to accuracy when making the decision about sharing, or the inability to distinguish between true and false content [6]. Individuals who are – when asked – able to distinguish between true and false content, but consciously or unconsciously do not prioritise content accuracy when sharing, have been found to do so for the following reasons. First, when sharing content on social media, individuals are prone to prioritising social considerations, such as how sharing certain content affects their social status. Users might also use the content they share as a way to connect to others [7, 8, 9]. Second, it has been found that the high volume of information that individuals are required to process on social media results in less attention dedicated to each piece of content that the individual comes across [10]. Finally, the inability to distinguish between true and false content has been shown to be associated with characteristics such as limited cognitive skills [11], lack of analytical reasoning [12], and lower levels of digital literacy [13, 14].
Interventions on social media have been shown to be successful across various contexts. Examples include prompting users to think about accuracy [11], and ‘inoculating’ users with small doses of misinformation and then refuting that piece of misinformation to increase awareness of how one can be manipulated into trusting content [15]. Given their success, it is compelling to explore ways in which social media users more likely to engage with misinformation sharing could be targeted with such interventions. While previous research presented above provides ample evidence on where to look for identifying individuals prone to misinformation sharing, obtaining data on these predictor characteristics can be complicated and/or expensive [16].
In this paper, we set out to provide a low-barrier approach to identify users that are more likely to share misinformation, relying on insights from previous literature, and easy-access data from X (Twitter). Our aim in utilising X metrics as accessible proxies for the above outlined phenomena is to demonstrate their effectiveness as initial predictors of users’ susceptibility to misinformation sharing. We aim to show that with the use of basic information available on X accounts, it is possible to identify users who are more likely to share misinformation, providing a basis for low-cost interventions in the future.
For this purpose, we selected four metrics to test. Our criteria for selecting these metrics were their easy accessibility and their ability to serve as proxies for the significant determinants of misinformation sharing found in the literature. Our metric selection was informed by the literature to ensure efficiency, however we prioritised data accessibility instead of the selection of perfect proxies. Based on these criteria, we chose the following metrics to represent three lines of misinformation sharing determinants: follower count for social considerations, average daily tweet count and number of users followed for cognitive overload, and account age for digital literacy.
Our hypothesis for each metric is that it could serve as a useful indicator for distinguishing between low and high factuality users, with specific expectations for how each metric would affect user factuality. We expect that users with more followers and the ones who tweet frequently and follow more accounts are less factual, while users with a longer life on social media are more factual. Specifically, based on the literature on social considerations shifting attention away from accuracy we expect that users with a higher number of followers would have lower factuality scores than those with a lower number of followers [6]. Furthermore, we expect to find that the higher the number of tweets one shares and the higher the number of accounts one follows (i.e. the higher the information exposure), the less the attention each piece of content would receive, as suggested by [10]. Finally, we expect to find the set of high factuality users to be dominated by users with older accounts, and the set of low factuality users with newer accounts, based on work by [13].
Results
We analysed data collected from X in 2022 (at that time still known as Twitter), consisting of 1,670,127 users and 14,688,374 tweets. Each user in the dataset is assigned a factuality score based on the credibility of the news sources they share, determined from their last 500 tweets containing a URL. Using these scores, we categorised users into high, middle, and low factuality groups (see Methods). Additionally, we retrieved various social network metrics on the same users, including follower count, number of followed accounts, tweet count, and registration date. Finally, we created 1,000 randomised versions of the dataset, where we shuffled users’ factuality scores while keeping all other features fixed, to serve as a basis for testing the statistical significance of our results (see Methods for more details on the dataset).
In the following, we present our findings on easy-access social network metrics’ effectiveness in the initial identification of individuals who tend to share a large amount of misinformation. We draw on the literature introduced above to guide the selection of the social network metrics tested. Specifically, drawing upon three strands of relevant literature – on social motivations, limited attention, and digital literacy – we select four metrics: follower count, daily average tweet count, followed account count, and the number of days since registration.
Popularity is associated with lower factuality
Building on the literature discussed above, highlighting the dominance of social considerations in social media sharing over accuracy concerns [6], we selected follower count on X (popularity) as our first metric. We hypothesise that follower count on X will serve as an indicator to identify users more inclined to share misinformation. To this end, we employed one-sided Mann-Whitney U (MWU) tests to compare the follower count distributions of the low and high factuality users, and verified their statistical significance by comparing them with the MWU statistics obtained on the 1,000 randomised datasets. Low factuality users were on average found to have a higher number of followers compared to high factuality ones, confirming our initial expectations (Figure 1a). The results were found to be statistically significant by comparing the MWU test statistic with the MWU test statistics obtained using the 1000 shuffled datasets (Figure 1b). This finding thus suggests that follower count can serve as an initial indicator in distinguishing between low and high factuality users, confirming our hypothesis.
Figure 1: Low factuality users tend to have higher follower count than high factuality users. Panel a: Comparing the distributions of follower count between low (orange curve) and high factuality (blue curve) users, we find that low factuality users have a significantly higher number of followers compared to high factuality users (p-value $<$ 0.0001). The median value for low factuality users (orange dotted line) is 1635 followers, whereas for high factuality users (blue dotted line) it is 1155. Panel b: The MWU test score obtained from the empirical data (red dotted line) is higher than all the MWU test scores calculated on the 1,000 shuffled datasets (green bars).
High tweeting rate and followed account count are also associated with lower factuality
Based on the related literature on information overload [10], we selected the number of tweets per day and the number of accounts followed as two additional metrics for distinguishing between low and high factuality users, and hypothesise that these two metrics are useful in distinguishing between low and high factuality users. Similarly to the previous metric, we utilised MWU scores to compare the distributions of these two metrics between low and high factuality users, and validated their significance against the 1,000 shuffled datasets. Low factuality users were observed both to post a higher number of tweets per day (Figure 2a) and to follow a greater number of users compared to their high factuality counterparts (Figure 3a). Both results were found to be statistically significant by comparing the MWU test statistics obtained from the empirical data to the MWU test statistics obtained from the 1000 shuffled datasets (see Figure 2b and Figure 3b, respectively). These results suggest that the two metrics tested can be useful in distinguishing between low and high factuality users.
Figure 2: Low factuality users tend to have higher tweet count than high factuality users. Panel a: Comparing the distributions of tweet count between low (orange curve) and high factuality (blue curve) users, we find that low factuality users have a significantly higher number of tweets compared to high factuality users (p-value $<$ 0.0001). The median value for low factuality users (orange dotted line) is 2.99 tweets per day, whereas for high factuality users (blue dotted line) it is 1.78. Panel b: The MWU test score obtained from the empirical data (red dotted line) is higher than all the MWU test scores calculated on the 1,000 shuffled datasets (green bars).
Figure 3: Low factuality users tend to have higher followed account count than high factuality users. Panel a: Comparing the distributions of followed account count between low (orange curve) and high factuality (blue curve) users, we find that low factuality users have a significantly higher number of followed accounts compared to high factuality users (p-value $<$ 0.0001). The median value for low factuality users (orange dotted line) is 2455 followed accounts, whereas for high factuality users (blue dotted line) it is 1914. The peak around 5000 is due to an X policy that limits the number of new followed accounts until the user obtains more followers. Panel b: The MWU test score obtained from the empirical data (red dotted line) is higher than all the MWU test scores calculated on the 1,000 shuffled datasets (green bars).
Longer social media presence is associated with higher factuality
Based on the literature on the relationship between digital literacy and misinformation sharing behaviour [13], we selected the number of days since registration as our last key metric for distinguishing between low and high factuality users. We employed MWU scores to compare distributions and ensured their statistical significance through comparisons with the 1,000 shuffled datasets. Low factuality users displayed a lower number of days since registration compared to high factuality users, i.e. older accounts tend to be more factual, aligning with our expectation (Figure 4a). The results were found to be statistically significant by comparing the MWU test statistics obtained from the empirical data to the MWU test statistics obtained from the 1000 shuffled datasets (Figure 4b). This finding supports the relevance of this metric in distinguishing between low and high factuality users.
Figure 4: Low factuality users tend to have lower number of days since registration than high factuality users. Panel a: Comparing the distributions of the number of days since registration between low (orange curve) and high factuality (blue curve) users, we find that low factuality users have a significantly lower number of days compared to high factuality users (p-value $<$ 0.0001). The median value for low factuality users (orange dotted line) is 3134 days since registration, whereas for high factuality users (blue dotted line) it is 3493. Panel b: The MWU test score obtained from the empirical data (red dotted line) is lower than all the MWU test scores calculated on the 1,000 shuffled datasets (green bars).
Combined effect of social network characteristics on factuality
Next, we employed multinomial regression to move from analysing the associations of factuality with individual social network features separately to considering the features all together. We do this in two stages. First, we perform multinomial regression with no interactions between the independent variables, second, we add interactions between explanatory variables. With conducting regression analysis, we are able to examine the effect of each key metric while simultaneously controlling for the effects of all other metrics. Thus, these models help us analyse the interplay of the metrics on factuality, and evaluate the extent to which these social network characteristics predict factuality scores.
We considered factuality as the outcome variable with three levels—low, middle, and high factuality. To capture the factors associated with notably low or high factuality, in each analysis we chose middle factuality to be the reference group that the other two are compared to. Explanatory variables included the four X metrics used above: follower count, followed account count, tweets per day, and days since registration. The explanatory variables were standardised for comparability. To measure the effects, we assessed the average marginal effects (AMEs) of the four social network metrics, as well as interactions between them. According to the results, each of the four social network metrics was found to have a significant relationship with both the low and high factuality groups—compared to the middle factuality group—with a number of significant interactions between variables.
First, we consider and report the AMEs of the four social network metrics on factuality (Figure 5). Examining the AMEs of follower count on factuality group membership, we find that for both low and high factuality groups, holding other variables constant, the more followers the user has, the more likely they are to be in these groups, compared to the middle factuality group. This effect is slightly but significantly stronger for the low factuality group than for the high factuality group. However, as the effect of followers points towards the same direction for both low and high factuality groups compared to middle factuality, this characteristic appears to be a weak metric for distinguishing between low and high factuality users, and is rather fit to distinguish low and high factuality users from middle factuality ones.
The number of users followed is negatively associated with both the high and low factuality groups, meaning that holding all other variables constant, following a high number of accounts makes a user more likely to be in the middle factuality category, while it makes them less likely to be in the low or high factuality categories. These findings are on the one hand consistent with our expectations, as we see a negative association between the number of users followed and high factuality, compared to middle factuality. Opposite to our expectation, we also see a negative association between the number of users followed and belonging to the low factuality category compared to belonging to the middle factuality category, however this effect is significantly weaker than the one belonging to high factuality. These effects show that there is a reverse U-shaped relationship between followed account count and the three factuality groups, both with low and high factuality users typically following fewer users than middle factuality users. These results suggest that the power of followed account count as a measure to classify low and high factuality users is therefore limited.
The average number of tweets produced by the user per day is negatively associated with being in the high factuality group compared to the middle factuality group, i.e. holding all other variables constant, high factuality users tweet less on average. For the low factuality group, we find a positive association compared to the middle factuality group, i.e. holding all other variables constant, low factuality users tweet more on average. These results are in line with our hypothesis and previous results, confirming that average daily tweet count is a useful indicator in distinguishing between low and high factuality users.
For days since registration, we see a positive effect for the high factuality, and the opposite for the low factuality group compared to the middle one. The results are therefore consistent with the outlined expectations and hypothesis, indicating that, holding all other variables constant, longer presence on X is associated with higher factuality, and that the metric is useful in distinguishing between low and high factuality users.
Figure 5: Average marginal effects of social network characteristics on factuality. All examined social network metrics are significantly associated with factuality, and the differences in the effects between low factuality (marked with orange) and high factuality (marked with blue) users is significant in all four cases. Zero effect would mean users are equally likely to be low or high factuality as middle factuality. Tweet count: higher tweet count means the user is more likely to be low factuality, and less likely to be high factuality than middle factual. Followed account count: higher followed account count means the user is less likely to be low factuality, as well as to be high factuality, but the latter effect is stronger. Follower count: higher follower count means the user is more likely to be low factuality, as well as to be high factuality, with the former effect being somewhat stronger. Days Since Registration: higher number of days means the user is more likely to be high factuality, and less likely to be low factuality.
Finally, we examined the interactions between the effects of the four social network metrics on factuality to unfold potential interplay of the independent variables. We have found that some of the social network metrics interact significantly with one another on factuality. To illustrate the interactions we divided the users into two equally sized groups based on the given social network metrics, e.g. follower count, and plotted the effects at the median of each group.
Our analysis reveals differences regarding the effect of the number of users followed on factuality (Figure 6, panel a). The effect of followed account count on high factuality is strongly negative among those with smaller follower base (less popular users), while slightly positive among those with larger follower base (popular users). In other words, less popular users are less likely to exhibit high factuality as they follow more users; whereas popular users show a slightly increased likelihood of high factuality with more followed accounts. This shows that the effect depends heavily on the number of followers the user has. Furthermore, this result shows that this effect is different from the main effect of followed account count, as the main effect showed a negative relationship, but in the case of popular users we found a slightly positive one. We can observe that among those who follow a low number of accounts, less popular users are more likely to be highly factual; while among those who follow many accounts, popular users are more likely to be highly factual. For the low factuality group similar trends are observed for both categories of follower count, where both less popular and popular users are less likely to exhibit low factuality as they increase the number of users they follow, with somewhat stronger effect for popular users.
In relation to average daily tweet count, our analysis (Figure 6, panel b) shows that as users increase their tweet activity, they are less likely to belong to the high factuality category and more likely to belong to the low factuality category, regardless of their number of followers. However, we do find differences in the strength of the effect. In the high factuality group the effect is more pronounced among less popular users, compared to the effect observed among popular users. Among those with a low number of average daily tweets, less popular users are more likely to be highly factual; while among those who tweet a lot, more popular users are more likely to be highly factual. Conversely, in the low factuality group, the relationship is less pronounced among less popular users. There we can observe that among those who only post a few tweets on average there is no difference between less popular and popular users; however, among those who tweet a high amount, popular users are more likely to belong to the low factuality group than less popular users.
In examining the interactions of the effects among the rest of the explanatory variables we found only slight differences. In none of the other cases was there a difference in the direction of the effect based on another variable. The observed differences are small, regarding either the magnitude or the strength of the effects (see the full set of explanatory variable interactions in Supplementary Figure 10).
Figure 6: Average marginal effect interactions between follower count and followed account count and average daily tweet count. Panel a: The effect of the number of accounts followed on factuality differs between users with few and many followers. The effect is opposite for the high factuality group, and different in strength for the low factuality group. Panel b: The effect of the average daily tweet count on factuality differs between users with few and many followers. The effect is opposite for the high factuality group, and different in strength for the low factuality group.
Discussion
In this study, we aimed to provide a low-barrier approach to distinguishing between individuals who are more and less likely to share misinformation on social media. Drawing upon existing literature and leveraging easily accessible data from X, we tested four social network metrics as potential predictors of user factuality. Our analysis focused on follower count, average daily tweet count, number of users followed, and account age. The results offer several insights into the effectiveness of these metrics as indicators of misinformation sharing tendencies.
Our findings indicate that when considered individually, each of the four metrics—follower count, average daily tweet count, number of users followed, and account age—has a significant relationship with user factuality scores. However, when all metrics were analysed together using multinomial regression, the distinctive power of follower count and followed account count diminished for the separation of low and high factuality, while average daily tweet count and number of users followed remained robust predictors of low and high factuality scores.
Higher average daily tweet counts were linked to lower factuality scores, consistent with our outlined expectation grounded in previous research which suggests that the sheer volume of content users engage with can detract from their ability to discern true from false information. This may be because the higher the number of tweets one shares, the less attention each individual piece of content would receive, making the user more likely to share non-accurate content [10].
Older accounts were associated with higher factuality scores, in line with our expectations that longer presence on the platform correlates with higher content accuracy, which may be moderated by better digital literacy. The finding might be explained by the fact that users with older accounts might have developed stronger digital literacy (specifically on X) as they have had more time to do so, whereas this is not yet true for users with newer accounts, in line with findings from [13]. Furthermore, according to diffusion of innovations theory [17], those who adopt new technologies earlier are often more digitally literate, which could explain why older accounts are more factual. We might also find fewer old accounts with low factuality because of certain policies of X removing strongly non-factual accounts.
Finally, the multinomial regression analysis revealed certain significant interaction effects in the case of follower and followed account count. Among those who follow a low number of accounts, less popular users are more likely to be highly factual, while among those who follow many accounts, popular users are more likely to be highly factual. Similarly, among those with a low number of average daily tweets, less popular users are more likely to be highly factual, while among those who tweet a lot, more popular users are more likely to be highly factual. These cross-effects highlight the context-dependent nature of the relationship between these social network metrics and factuality.
These findings have several important implications. For social media platforms, the results suggest that basic metrics like tweet count and account age could be used to identify users more likely to share misinformation, allowing for targeted interventions such as accuracy prompts or digital literacy campaigns. Policymakers could advocate for the incorporation of these metrics into misinformation mitigation strategies, ensuring a low-cost, scalable approach. Future research should continue to explore additional accessible data points and refine existing metrics to enhance the accuracy of identifying misinformation-prone users.
In conclusion, this study demonstrates that basic, easily accessible social network metrics, particularly tweet count and account age, can effectively differentiate between low and high factuality users. These studied metrics provide a practical and efficient method for identifying users susceptible to sharing misinformation, offering a foundation for cost-effective interventions to curb the spread of misinformation on social media platforms.
Methods
Data
Our study builds upon prior studies by Rathje et al. [18] and (Galeazzi et al. 2023), and to carry out the analysis we rely on the X (Twitter at the time) dataset collected by these authors. Rathje et al. [18] gathered 463 survey responses through the online recruitment platform Prolific. Participants were asked to respond to survey questions and requested to provide their X handles. Building on this, Galeazzi et al. (2023) obtained the usernames of all those Prolific participants who provided their X handles, then collected the X handles of each Prolific participant’s followers, resulting in a dataset of a total of n = 1,670,127 users. Finally, the authors collected the latest 500 tweets of each user that contains an URL to external sources, covering a period up to a year before the date of their latest tweet, producing a total of n = 14,688,374 X tweets. From this dataset we obtained all users with available factuality scores, which resulted in a dataset of 19,596 users. Furthermore, we utilise data on the number of followers of the user, number of people followed by the user, total number of tweets produced by the user, registration dates, and other details, all directly extracted from X.
User factuality
To measure the credibility of news outlets linked in X tweets, Galeazzi et al. (2023) introduce a metric, referred to here, and in their paper, as Factuality. The quantification of this metric relies on data from Media Bias Fact Check (MBFC), an independent fact-checking organisation that classifies news outlets based on both their reliability and political bias [19]. Each news outlet on MBFC is assigned labels indicating its reliability (Very Low, Low, Mixed, Mostly Factual, High, Very High). In the context of their study, news outlets’ factuality corresponds to MBFC’s reliability label. These metrics are applied individually to each X user (denoted as u), generating a vector, $F_{u}\in\mathbb{R}^{6\times 1}$, representing the percentage of links shared by the user for Factuality categories. In this paper we rely on the average of the scores related to the tweets posted by the same user to assign a unique factuality score to each user, which will serve as the basis of our analysis. We segmented users into low, middle, and high factuality groups based on their respective factuality scores. This involved dividing users into the three factuality categories such that the bottom and top 25%, 30%, or 35% would be the low and high factuality users, respectively, and the remaining middle would be the middle factuality group, resulting in three distinct factuality groups. While the main focus of our paper lies on the results derived from the factuality categorisation taking the top and bottom 30% as low and high factuality users, respectively, we also present findings obtained applying the other two thresholds in the Supplementary Information to validate the robustness of our conclusions.
Filtering the dataset for regular users
As the purpose of our study is understanding the behaviour of “human users” [20], we initially curated our datasets to focus on users who are more representative of organic human social media users. I.e., we wanted to exclude accounts from the study that appeared to be either bots, business, political or celebrity accounts. To ensure the selection of regular individuals, we implemented specific criteria as outlined below. These criteria were derived from various sources and were applied iteratively to curate our dataset.
Verified status: Verified accounts, typically belonging to recognized individuals or brands, were identified and removed from the dataset as they do not represent typical user behaviour. This was based on X’s verification policy—a rigorous review process of the authenticity, notability, and activity of the account at the time of the data collection [21]. Such accounts comprise 2.9% of all accounts in the dataset.
Average number of tweets per day: Accounts with an exceptionally high average number of tweets per day were considered unlikely to belong to regular users, as this behaviour is more characteristic of commercial or political accounts, sometimes bots [22]. We set thresholds to remove accounts with tweet/day ratios exceeding certain values expected to be limits of regular user behaviour.
Number of followers: Accounts with a very high number of followers were considered atypical for regular users, as such levels of fame are usually attained outside of X or through exceptionally high activity on the platform. We removed accounts with follower counts surpassing a certain threshold.
Follower/followed account ratio: A high proportion of followers compared to followed accounts is characteristic of celebrity-like accounts rather than typical users. Therefore, we removed accounts with follower/followed account ratios exceeding certain values.
Number of followed accounts: Accounts with an unusually high number of followed accounts may indicate bot-like behaviour, particularly if the followings are not reciprocal [23]. We set thresholds to remove accounts with followed account counts exceeding specific values.
Followed account/follower Ratio: A high proportion of non-reciprocal followings may also indicate bot-like behaviour [23]. Therefore, accounts with followed account/follower ratios exceeding certain values were removed. Based on the stringency of the above introduced criteria we created three distinct groups of datasets: relaxed, middle, and strict. By applying these criteria, we aimed to ensure the selection of regular individuals whose behaviour is more representative of typical social media users for our analysis. The specific values for each strictness category are shown in Table 1.
Criteria Relaxed Middle Strict
Verified Status - Not Verified Not Verified
Average Tweets Per Day $\leq 56$ $\leq 32$ $\leq 16$
Follower Count $\leq 10,000$ $\leq 10,000$ $\leq 5,000$
Follower/Followed Account Ratio $\leq 10$ $\leq 5$ $\leq 3$
Followed Account Count $\leq 10,000$ $\leq 10,000$ $\leq 5,000$
Followed Account/Follower Ratio $\leq 10$ $\leq 10$ $\leq 5$
Table 1: Criteria for dataset variation for robustness check.
As in the case of factuality grouping, the main focus of our paper lies on the results derived from the middle category, but we also present findings obtained applying the two other criteria in the Supplementary Information to validate the robustness of our conclusions. The number of users in each dataset are shown in Table 2.
Criteria Relaxed Middle Strict
Top/bottom 25% 4899 4612 3172
Top/bottom 30% 5878 5535 3807
Top/bottom 35% 6858 6457 4441
Total 19596 18450 12690
Table 2: Number of observations in each of the datasets used for validation of results.
Randomised dataset
To ensure the conclusions drawn from our results are not due to chance, we constructed a counterfactual dataset consisting of 1,000 datasets mirroring the structure of our original dataset. In these counterfactual datasets, factuality scores were randomly shuffled across individuals while keeping all other characteristics constant.
Analysis of combined effects
We employed multinomial regression analysis to investigate the combined effects of our four social network metrics—follower count, followed account count, tweets per day, and days since registration—on factuality, which is categorised into three levels: low, middle, and high. Multinomial regression is particularly suitable for this analysis as it allows for the modelling of outcomes with multiple discrete categories, enabling us to examine the influence of multiple predictors on the probability of an observation belonging to each factuality level simultaneously. To interpret the results, we utilised Average Marginal Effects (AMEs). AMEs offer a clear and intuitive measure of the impact of each social network metric on the probability of an account being classified into each factuality category. By translating regression coefficients into average changes in probabilities, AMEs enhance the interpretability of our findings, making it easier to understand and compare the effects of different metrics. This approach aligns with the recommendations by [24], who highlights the advantages of using marginal effects in complex regression models.
Acknowledgements
J.K. acknowledges funding from the Hungarian Academy of Sciences Lendület Program: LP2022-10/2022.
Author contributions statement
All authors designed the study. A.G. collected and processed the data. J.S. carried out the data analysis and statistical modeling. E.O., J.K. and A. G. supervised the study. All authors reviewed and approved the final manuscript.
Additional information
Data availability. The X (Twitter) data is made available in accordance with X’s terms of service. Tweet IDs for each COP are available at: [the link will be provided upon publication]. The corresponding tweets can be downloaded using the official Twitter API (https://developer.twitter.com/en/docs/twitter-api).
Competing interests. The authors declare no competing interests.
Supplementary Information
Robustness checks
Distributions of follower count
Figure 1: Low factuality users tend to have higher follower count than high factuality users. Comparing the distributions of follower count between low (orange curve) and high factuality (blue curve) users, we find that low factuality users have a significantly higher number of followers compared to high factuality users (p-value $<$ 0.0001) in each of the tested datasets. The median value for low factuality users (orange dotted line) is higher than for high factuality users (blue dotted line) in each of the tested datasets.
Figure 2: Low factuality users tend to have higher follower count than high factuality users. The MWU test score obtained from the empirical data (red dotted line) is higher than all the MWU test scores calculated on the 1,000 shuffled datasets (green bars) in each of the tested datasets.
Distributions of tweet count
Figure 3: Low factuality users tend to have higher tweet count than high factuality users. Comparing the distributions of tweet count between low (orange curve) and high factuality (blue curve) users, we find that low factuality users have a significantly higher number of tweets compared to high factuality users (p-value $<$ 0.0001) in each of the tested datasets. The median value for low factuality users (orange dotted line) is higher than for high factuality users (blue dotted line) in each of the tested datasets.
Figure 4: Low factuality users tend to have higher tweet count than high factuality users. Panel b: The MWU test score obtained from the empirical data (red dotted line) is higher than all the MWU test scores calculated on the 1,000 shuffled datasets (green bars) in each of the tested datasets.
Distributions of followed account count
Figure 5: Low factuality users tend to have higher followed account count than high factuality users. Comparing the distributions of followed account count between low (orange curve) and high factuality (blue curve) users, we find that low factuality users have a significantly higher number of followed accounts compared to high factuality users (p-value $<$ 0.0001) in each of the tested datasets. The median value for low factuality users (orange dotted line) is higher than for high factuality users (blue dotted line) in each of the tested datasets. The peak around 5000 is due to an X policy that limits the number of new followed accounts until the user obtains more followers.
Figure 6: Low factuality users tend to have higher followed account count than high factuality users. The MWU test score obtained from the empirical data (red dotted line) is higher than all the MWU test scores calculated on the 1,000 shuffled datasets (green bars) in each of the tested datasets.
Distributions of account age
Figure 7: Low factuality users tend to have lower number of days since registration than high factuality users. Comparing the distributions of the number of days since registration between low (orange curve) and high factuality (blue curve) users, we find that low factuality users have a significantly lower number of days compared to high factuality users (p-value $<$ 0.0001) in each of the tested datasets. The median value for low factuality users (orange dotted line) is lower than for high factuality users (blue dotted line) in each of the tested datasets.
Figure 8: Low factuality users tend to have lower number of days since registration than high factuality users. The MWU test score obtained from the empirical data (red dotted line) is lower than all the MWU test scores calculated on the 1,000 shuffled datasets (green bars) in each of the tested datasets.
Regression results
Figure 9: Average marginal effects of social network characteristics on factuality for the 9 datasets for checking robustness of results.
Interactions
Through plotting the interactions of follower count with all other metrics, certain differences in the relationships between these metrics and factuality levels across users with varying follower counts have been observed. Specifically, our analysis reveals that in relation to followed account count, users with fewer followers are less likely to exhibit high factuality as they follow more users, whereas those with a larger follower base show a slightly increased likelihood of high factuality with more followed accounts (Figure 10, panel a).Thus, the effect of followed account count on high factuality is strongly negative among those with smaller follower base, while slightly positive among those with larger follower base. For the low factuality group similar trends are observed for both categories of follower count, where users with both few and many followers are less likely to exhibit low factuality as they increase the number of users they follow. Nevertheless, the effect is less strong among those with fewer followers.
In relation to average daily tweet count our analysis (Figure 10, panel b) shows that as users increase their tweet activity, they are less likely to belong to the high factuality category and more likely to belong to the low factuality category, regardless of their follower count. Nevertheless, among users with fewer followers, the effect is more pronounced in the high factuality group compared to the effect observed among users with a larger number of followers. Conversely, in the low factuality group, this relationship is less pronounced among users with fewer followers.
Regarding the relationship between factuality and days since registration, similar patterns emerge across different follower counts (Figure 10, panel c). We show that as users spend more time registered on the platform, they are more likely to belong to the high factuality category and less likely to belong to the low factuality category, and this is consistent in direction and magnitude across users, regardless of their follower count. However, the positive effect of days since registration on high factuality is stronger among those with lower number of followers compared to those with larger number of followers. Nevertheless, we cannot observe such difference in the effects on low factuality.
In examining average daily tweet count across different levels of followed account count and its association with factuality levels (Figure 10, panel d) we find that an increase in tweet activity is associated with a decreased likelihood of belonging to the high factuality group, regardless of followed account count, but the relationship is stronger among those who follow fewer accounts. In contrast, the relationship between tweet count and low factuality varies based on followed account count with a higher degree. Users, who follow a larger number of accounts demonstrate a decreased likelihood of belonging to the low factuality group as tweeting activity increases. Conversely, among users with who follow fewer accounts, higher tweet counts are associated with a somewhat increased likelihood of belonging to the low factuality group.
Our findings regarding the interaction between days since registration and followed account count (Figure 10, panel e) indicate that regardless of followed account count, users who have been registered on the platform for a longer duration are more likely to be in the high factuality category. Conversely, the relationship between days since registration and low factuality varies based on followed account count. The relationship is consistently negative between days since registration and low factuality, indicating that longer registration periods are associated with a decreased likelihood of low factuality, however, this effect is somewhat more pronounced among users who follow less accounts.
Examining the interaction between days since registration and tweet count in relation to factuality levels reveals consistent trends across users with different average daily tweet counts (Figure 10, panel f). Our analysis demonstrates that as the registration period lengthens, users are more likely to exhibit higher factuality levels, and less likely to exhibit lower factuality levels, irrespective of their tweet activity.
Figure 10: Interactions between various social network metrics in relation to belonging to low and high factuality groups.
Prediction Accuracy
Accuracy scores obtained from shuffled datasets against those derived from the real data to demonstrate that the accuracy obtained from our regression exceeds that expected by random chance.
Figure 11: Accuracy score of predictions using the four social network metrics on the real dataset (red dotted line) compared to the accuracy scores obtained on the 1000 randomised datasets (green bars).